OpenStudy (michael014):

Urgent help please! use the digits 1,3,5 and 7 once each, to complete the multiplication and get the smallest answer

OpenStudy (javk):

what do you mean by 'the multiplication', is it like a specific equation, or something?

OpenStudy (anonymous):

You kinda need to be more specific. Any order would give you the same answer.

OpenStudy (michael014):

Like this , there are four boxes, underneath and on top. I need to put the digits in there to get the smallest multiplication possible. I.e: two digits on top and two under neath with a multiplication X in middle. I need smallest possible answer.

OpenStudy (javk):

If you mean like this... Then you have to: a) Place the two smallest numbers in the tens column b) From the two remaining numbers, take the smaller of the two and place it in the units column, next to the smaller of the two that were placed in the tens column c) take the last remaining number and slot that in too you will end up with \[37\times 15\] Now go back and check \[13\times 57=741\] \[35\times 17=595\] \[37\times 15=555\] so your answer was correct